# Sales Whisperer: A Human-Inconspicuous Attack on LLM Brand Recommendations

### v1

@@ 7.1.4 Statistical analysis
For each of the six base and perturbed prompt pairs, we analyzed the difference between base and perturbed group responses with a series of non-parametric tests on our 11 main measurement variables: two-tailed Mann-Whitney U tests for Likert data and chi-squared tests for binary data. To understand whether the base and perturbed groups are equivalent, we replaced non-significant tests for difference with tests of equivalence. To establish equivalence, we used the two one-sided tests procedure (TOST) and set our equivalence margin to be for 80% power [142].666Testing for equivalence is a deviation from the pre-registration; however, we believe this approach paints a more complete picture. Open-ended brand-recall questions (e.g., what brand participants would pick based on the response), were coded into binary categories: whether the participant reported the targeted brand or not. hesitant or otherwise unclear responses did not count as a match.
@@ 7.2 Results
We first describe our participants and then present results. We find that, under most measures, perturbed prompts and corresponding responses are human-inconspicuous. Further, these prompts, in most measures, successfully nudge more users into noticing the target brand.
@@ 7.2.2 Statistical Evaluation
In order to measure attack success, we recorded the percentage of participants who would pick the targeted brand given the response, the percentage of participants who noticed the brand in the response, and the percentage of participants who said the targeted brand was the top recommendation. As shown in (Table III), in five out of six categories our attacks were able to increase the prominence of the target brand in at least one of the three questions: in four categories participants were more likely to pick the targeted brand when given the perturbed prompts, in five categories participants were more likely to notice the targeted brand, and in three categories participants were more likely to say the targeted brand was the top recommendation.
8 Concluding Discussion
We identify a novel threat model in which an adversary aims to induce target biases in LLM responses users receive by using adversarially crafted prompts. We specifically investigate whether adversaries can bias responses towards a target brand through prompt suggestions (e.g., recommended prompts within a chatbot, prompt sharing services).
We contribute a dataset of 449 prompts that ask for product recommendations over 77 product categories, and develop a method to evaluate which brands are mentioned in LLM responses using target words from the same dataset. We evaluate the difference in probabilities that brands are mentioned across similar prompts (paraphrases) within each product category. We find up to a difference between pairs of prompts. Next, we develop a method to perturb a base prompt by replacing words with their synonyms from our new synonym dictionary. We select a candidate using a loss function that correlates with measured attack success. This results in human-inconspicuous prompts that increase the probability that a target brand is mentioned. Novel with respect to other work that studies inconspicuous text manipulations, we perform a user study that demonstrates that, compared to base prompts, our synonym-replacement prompts are successful in promotinga target brand while remaining inconspicuous to humans.
We also explore transfer attacks of our synonym-replacement approach from open-sourced models where logits are available to GPT models, finding transferability between Llama3-it and GPT-3.5-Turbo.
Taken together, our work highlights the potential for adversarial prompt recommendations to bias LLM responses. While our focus is on brand recommendations, our methods are general and can be applied to other adversarial goals, such as propaganda, misinformation, and bias on social issues. As such, we suggest that defensive measures should be taken. Since the perturbations we introduce are imperceptible to humans, we suggest that users should be warned to check sources before using prompts they did not author. Notably, warnings about LLM use are already commonplace in many chatbot applications, ChatGPT noting “ChatGPT can make mistakes. Check important info.” [146]. We suggest that these be implemented for untrusted prompts.
@@ 8 Concluding Discussion
Throughout our experiments evaluating rephrasings in §6.1, we gathered completions for prompts on categories with products manufactured by Meta, Google, and Microsoft, which allowed us to examine how large language models developed by these companies perform when asked about product categories that include products manufactured by them. We evaluated the average score, as defined in §5.1, of all brands over all prompts for categories where one of the brands was Meta, Google, or Microsoft. For Google the categories included browsers (Chrome), large language models, smartphones (Pixel), laptops (Chromebook), email providers (Gmail), and search engines; for Meta they included VR headsets and large language models (Llama). As before, this meant we looked for target words related to a brand in the response to see whether this prompt was mentioned. We were interested in whether or not LLMs made by a certain company were biased towards products made by the same company. All results are shown in Fig. IV

### v2

@@ 5.1.4 Statistical analysis
For each of the six base and perturbed prompt pairs, we analyzed the difference between base and perturbed group responses with a series of non-parametric tests on our 11 main measurement variables: two-tailed Mann-Whitney U tests for Likert data and chi-squared tests for binary data. To understand whether the base and perturbed groups are equivalent, we replaced non-significant tests for difference with tests of equivalence. To establish equivalence, we used the two one-sided tests procedure (TOST) and set our equivalence margin to be for 80% power [118].777Testing for equivalence is a deviation from the pre-registration; however, we believe this approach paints a more complete picture. Open-ended brand-recall questions (e.g., what brand participants would pick based on the response), were coded into binary categories: whether the participant reported the targeted brand or not. Hesitant or otherwise unclear responses did not count as a match.
@@ 5.2 Results
We first describe our participants and then present the results. We find that, under most measures, perturbed prompts and corresponding responses are inconspicuous. Further, these prompts, in most measures, successfully nudge more users into noticing the target brand.
@@ 5.2.2 Statistical Evaluation
In order to measure attack success, we recorded the percentage of participants who would pick the targeted brand given the response, the percentage of participants who noticed the brand in the response, and the percentage of participants who said the targeted brand was the top recommendation. As shown in Table III , in five out of six categories our attacks were able to increase the prominence of the target brand in at least one of the three questions: in four categories participants were more likely to pick the targeted brand when given the perturbed prompts, in five categories participants were more likely to notice the targeted brand, and in three categories participants were more likely to say the targeted brand was the top recommendation.
Effect of earlier brand appearance
When generating perturbed prompts through synonym replacement, we assumed that earlier mention of a target concept would be more likely to influence users. Thus, we used a loss function that increases the probability of a target concept being mentioned at the beginning of the response. Here, we run an exploratory (not pre-registered) analysis to test this assumption.
Specifically, we ran a series of logistic regressions to predict whether participants would notice the target brands based on how soon the target brand was mentioned in the response. For formulated the model where correctly noticing the target brand is the dependent variable and the position of the target brand in the response is the independent variable (if “Verizon” is the 15th word, the position would be 15). We included a random effect for the prompt used, since product categories and base prompt phrasing might have different baseline effects. We find that, later mentions of target brands in responses are less likely to be noticed by participants. This observation remains true for all three measurements of attack success: what brand is picked, what the top perceived recommendation is, and if the brand is perceived as recommended by the llm. A 10-word increase in the position of the target brand reduces the chances of users noticing the target brand by 1.7%, 5.9%, and 0.6% respectively (all ).
6 Discussion
We identify a novel threat model in which an adversary aims to induce target biases in LLM responses users receive by using adversarially crafted but innocuous-seeming prompts. We further empirically show that users do not notice that these prompts are adversarial but are influenced by the biased responses that the prompts result in. Here we discuss the implications of our findings and suggest potential defenses.
Implications
This risk we highlight in this work stems from real-world deployments of LLMs in chatbots and the accompanying prompt providers. As such, it has real-world implications. Users may be manipulated into thinking a certain way, while being under the impression that they are receiving unbiased advice. Prompt providers could be giving the impression of a personalized experience, while in reality undermining the user’s autonomy. This risk can fundamentally be thought of as a risk of intentional algorithmic bias, which is well understood to be particularly dangerous in the context of political and social issues [122], but also concerning in commercial contexts [123].
6.1 Defenses: multi-pronged
We identify a series of defenses—with various tradeoffs between effectiveness and deployment cost—that different stakeholders can employ to mitigate the risk of adversarial prompts. The suggested defenses are complementary, and can, and we believe should be, used in conjunction.
User education and warnings
Using prompts from (untrusted) sources is akin to running untrusted code copied from untrusted sources, a well-studied problem in computer security [124, 125] . Thus, similar protection mechanisms, like warnings [126], might be effective. Unlike untrusted code though, inconspicuous attacks are difficult to detect and therefore might require more invasive warnings. Security warnings have a long history in human-centered design [127, 128, 129, 130], they can be integrated into the user interface to possibly help users make more informed decisions. These warnings could appear in various locations depending on the the owner of the prompt provider. If the prompt provider is not the chatbot owner, the chatbot could warn users about the potential risks of using prompts from untrusted sources. Notably, certain chatbots already warn users – ChatGPT warning, “ChatGPT can make mistakes. Check important info.” [131]. Further, prompt libraries can warn users about the potential risks of using prompts from other users, similar to proposed warnings in programming forums [126].
Warnings, however, are likely only part of the solution. More involved user education campaigns on how to use LLMs and chatbots safely and effectively might need to be developed [4].
Robust models
Our attacks fundamentally rely on LLMs (like other deep neural networks) being brittle, a fact we also observe in our work (§4.3.1). They respond to small changes in prompts in a way that is unexpected and can be manipulated. Despite the claims of robustness by many LLM providers, our work (and many others) show that there is still much work to be done. Existing robust LLMs focused on the correctness of LLM answers [132, 133, 134, 135, 136, 137], and to the best of our knowledge, we are the first to reveal that slight (inconspicuous) perturbation in LLMs can lead to vast differences in LLMs’ likelihood of mentioning concepts. Such differences may lead to bias and risks that hurt user autonomy, including but not limited to advertising without user awareness and propaganda, as we showed in this paper. We suggest LLM providers emphasize robustness against these defenses. However, given the rapid—perhaps rushed—deployment of LLMs in the wild (e.g., [138]), it might be up to regulators to enforce robustness requirements [139].
New Bias Metrics
Our work does not use prior definitions of bias [47, 48] during evaluation. This is intentional: existing bias metrics focus on discrimination, hate speech, and exclusion, but do not capture the type of bias we introduce in this work. For instance, many brands could be mentioned in an LLM response, and an ideal bias metric should be able to capture how much each one can be biased. This suggests that new metrics need to be developed to capture this type of bias, enrich the definition of LLM robustness, and eventually, make testing meaningful.
Continous audits
In the absence of guaranteed robustness, an elusive target, frequent testing of models for bias and other risks is essential. We advocate for systematic testing of prompt providers for bias, including the type of bias we introduce in this work. Such a system could regularly check if prompts are biased towards certain concepts, and if so, alert users. The effectiveness of such a system would fundamentally depend on the quality of the bias detection tests. This brings us to our next suggestion.
We further find that running measurements on LLMs is difficult due to the variability in responses. While this is a feature that LLM developers intentionally build, it also complicates the measurement of bias. How can we measure the underlying bias in the model? Is simply averaging a large number of responses enough? We advocate for future research to explore this question.
7 Conclusion
We identify a novel threat model in which adversaries bias LLM responses in a target direction by suggesting subtly altered prompts to unsuspecting users. Through a series of experiments and a user study, we develop this attack and demonstrate that we can bias LLM responses towards a target concept while remaining inconspicuous to humans. We further show that this attack can be used to bias LLMs in harmful and helpful ways, and that the attack is transferable between models. These findings highlight the risk of adversaries introducing bias in LLM responses in a unique way, suggesting a need for defensive measures like warnings and robustness requirements.
@@ 7 Conclusion
Throughout our experiments evaluating rephrasings in §4.3.1, we gathered completions for prompts on categories with products manufactured by Meta, Google, and Microsoft, which allowed us to examine how large language models developed by these companies perform when asked about product categories that include products manufactured by them. We evaluated the average score, as defined in §4.2.1, of all brands over all prompts for categories where one of the brands was Meta, Google, or Microsoft. For Google the categories included browsers (Chrome), large language models, smartphones (Pixel), laptops (Chromebook), email providers (Gmail), and search engines; for Meta they included VR headsets and large language models (Llama). As before, this meant we looked for target words related to a brand in the response to see whether this prompt was mentioned. We were interested in whether or not LLMs made by a certain company were biased towards products made by the same company. All results are shown in Fig. IV